OpenStudy (anonymous):

suppose an investment of 10,000 doubles every 13 years. How much is the investment worth after 65 years?

OpenStudy (anonymous):

\[10000\times2^\frac{65}{13}\]

OpenStudy (anonymous):

\[\frac{65}{13}=5\] \[2^5=32\] \[10000\times 32=320000\]